A distribution of measurements is relatively mound-shaped with a mean of 80 and a standard deviation of 14. Use this information to find the proportion of measurements in the given interval. Greater than 94
Answer #1

Given,

= 80 , = 14

We convert this to standard normal as

P(X < x) = P(Z < x - / )

So,

P(X > 94) = P(Z > 94 - 80 / 14)

= P(Z > 1 )

= 1 - P(Z < 1)

= 1 - 0.8413

= 0.1587